OpenStudy (blutz):

if the sides of a triangle are 16,20 and 27 respectively what is the length of the bisector of the largest angle

KjSaif (kjsaif):

should there be a picture?

KjSaif (kjsaif):

you there @blutz

OpenStudy (eliesaab):

Use the formula on https://proofwiki.org/wiki/Length_of_Angle_Bisector

OpenStudy (eliesaab):

\[d^2=\frac{b c \left((b+c)^2-a^2\right)}{(b+c)^2}\]

OpenStudy (eliesaab):

Replacing and doing some arithmetic gives \[ d^2 = 140\\ d= 2 \sqrt{35} \]

OpenStudy (eliesaab):

@KJSaif

KjSaif (kjsaif):

he left @eliesaab

KjSaif (kjsaif):

@blutz left

OpenStudy (eliesaab):

OK